Square root y+10=y-2
Answer by checkley79(3341)

Square root y+10=y-2 SQUARE BOTH SIDES.
Y+10=Y^2-4Y+4
Y^2-4Y-Y+4-10=0
Y^2-5Y-6=0
(Y-6)(Y+1)=0
Y-6=0
Y=6 ANS.
Y+1=0
Y=-1 ANS.